OpenStudy (anonymous):

PLEASE HELP :( Solve 3x^2+x+10=0

OpenStudy (cwrw238):

this won't factor so you need to use the quadratic formula for ax^2 + bx + c = 0 x = [-b +- sqrt(b^2 - 4ac) ] / 2a

OpenStudy (cwrw238):

here a = 3, b = 1 and c = 10

OpenStudy (anonymous):

Thank you I'm gonna try holdon

OpenStudy (anonymous):

@cwrw238 i go -1 sq -119/6

OpenStudy (cwrw238):

= [-1 + sqrt(1 - 4*3*10)] / 6 =[ -1 + sqrt(-119) ] = 1/6 + ( 1/6) sqrt119 i or 1/6 - (1/6) sqrt119 i these are 2 complex numbers ( i = sqrt(-1) )

OpenStudy (anonymous):

@cwrw238 here are the answer choices. -2.83 and .83 No real solutions -2.01 and 1.67 -1.67 and 2.01

OpenStudy (cwrw238):

ok - complex numbers are partly real and partly imaginary ( thats the part with 'i' in it) so the correct choice is No real solutions